OpenStudy (anonymous):

The position of an object at time t is given by s(t) = 1 - 10t. Find the instantaneous velocity at t = 10 by finding the derivative.

OpenStudy (anonymous):

Do you know how to find derivatives?

OpenStudy (anonymous):

@Data_LG2 kind of.... i'm like really confused on the whole process

OpenStudy (anonymous):

can you tell me the derivative of the function? s(t) = 1 - 10t

OpenStudy (anonymous):

-9????? @Data_LG2

OpenStudy (anonymous):

i could be wrong..

OpenStudy (anonymous):

nope okay... let's start from the basic remember that the derivative of any real number is zero for the function s(t) = 1 - 10t , one is a whole number.. next, the product rule \(f(x)=x^n\) -> \(f'(x)=nx^{n-1}\) for the function you have -10 t , the derivative of this will be -10 only therefore , \(s(t)=1-10t\) --> \(s'(t)=-10\) now find \(s'(10)=?\) it will be

OpenStudy (anonymous):

just 10?

OpenStudy (anonymous):

don't forget the negative :)